In a Young’s double slit experiment, the two slits act as coherent sources of waves of equal ampiitude A and wavelength λ. In another experiment with the same arrangement the two slits are made to act as incoherent sources of waves of same amplitude and wavelength. If the intensity at the middle point of the screen in the first case is I1 and In the second case is I2 then the ratio I1/I2 is

Grade:

1 Answers

Arun
6 years ago
Dear Sarthak
I coherent = n² * Io = 2² * Io = 4 Io
I incoherent = Io + Io = 2 Io
Hence ratio = 4/2 = 2/1
Regards
Arun (askIITians forum expert)